Calculate the density of neon gas ( in g/L) at 758 mmHg and 52.3°C.
Answer #1

According to the ideal gas law,

PV = nRT

=> V= nRT/P

= (W/M) RT/P

=> v= (WRT/PM)

=> PM/ RT = (W/V)

=> d = (PM/RT)

Where, P= pressure in atm V= Volume in liter, R= universal gas constant in lit atm mol -1 K -1 T= temperature at K ,

M = molecular weight, W= weight in grames, and W/V = d , is density in gm/ L

Now, pressure (P) = 758/760 = 0.9974 atm

T = (52.3 + 273.15) = 325.45 K

Molecular weight of Neon (M) = 20.1797 g/mol

So, density = { (0.9974× 20.1797) ÷ ( 0.082 × 325.45)}

= 0.737 g/L